### v4

# Collaborative Causal Sensemaking: Closing the Complementarity Gap in Human–AI Decision Support

Abstract.
LLM-based agents are increasingly deployed for expert decision support, yet human-AI teams in high-stakes settings do not yet reliably outperform the best individual. We argue this complementarity gap reflects a fundamental mismatch: current agents are trained as answer engines, not as partners in the collaborative sensemaking through which experts actually make decisions. Sensemaking (the ability to co-construct causal explanations, surface uncertainties, and adapt goals) is the key capability that current training pipelines do not explicitly develop or evaluate. We propose Collaborative Causal Sensemaking (CCS) as a research agenda to develop this capability from the ground up, spanning new training environments that reward collaborative thinking, representations for shared human-AI mental models, and evaluation centred on trust and complementarity. Taken together, these directions shift MAS research from building oracle-like answer engines to cultivating AI teammates that co-reason with their human partners over the causal structure of shared decisions, advancing the design of effective human–AI teams.
Key words and phrases:
Human-AI Collaboration, Multi-Agent Systems, Collaborative Sensemaking, Causal Reasoning, Human–AI Complementarity, Trust
1. Introduction
Multi-agent systems (MAS) built from large language model (LLM) agents are increasingly positioned as decision-support teammates for humans in domains such as personalisation, planning, and multi-objective optimisation, where consequences are delayed, uncertain, and value-laden marivate2014quantifying; roijers2013survey; kaelbling1998planning; sutton2018reinforcement; wang2019bayesian. While AI assistants have unlocked productivity gains in verifiable domains like coding and translation, empirical work in decision-making under uncertainty reveals a persistent complementarity gap: where judgement is subjective and verification is costly, human–AI teams frequently underperform the best individual agent bucinca2024towards; hemmer2024complementarity; fragiadakis2025evaluating; steyvers2022bayesian; rastogi2022taxonomy. For next-generation MAS, this is not a minor usability flaw but a core systems failure: agents that cannot sustain calibrated, shared understanding with their human partners will systematically mis-coordinate, even if their standalone predictions are strong.
A growing body of studies documents characteristic failure modes that undermine calibrated trust. Users over-weight confident model outputs even when these conflict with domain expertise, exhibiting automation bias and over-reliance goddard2012automation; lyell2017automation; alonbarkat2023trust; Bansal2021UpdatesHarm. Verification-and-correction loops can erase efficiency gains, as experts feel compelled to second-guess model suggestions step by step Bansal2021UpdatesHarm; bucinca2024towards; hemmer2024complementarity. Alignment methods that reward agreement and user satisfaction can induce sycophancy, where models collapse to the user’s prior beliefs even when these conflict with evidence perez-etal-2023-discovering; sharma2024towards. This is fatal for sensemaking, which by definition requires the repair and restructuring of mental models, not merely their confirmation weick1995sensemaking; klein2006dataframe. The result is trust poorly calibrated to actual competence: humans rely on agents for fluency rather than causal reasoning zhang2020appropriatereliance; gomez2025taxonomy; lai2023designspace.
Current training pipelines do not address this. Preference-based alignment (RLHF, DPO, and variants) shapes outputs toward helpfulness and safety ouyang2022instructgpt; awesome_llm_posttraining; rafailov2023dpo; dvo; cpo; reasoning methods (chain-of-thought, RL with verifiable rewards, process supervision) make multi-step reasoning instrumentally useful wei2022chainofthought; wei2022emergent; havrilla2024rlvr; luong2024reft; zelikman2024quietstar; and world-model approaches train predictive models of environment dynamics hafner2020dreamer; hao2023llmreasoners. However, these methods optimise for solitary performance: they align the agent to a label, a verifier, or a simulator, not to the evolving mental model of a partner. Any collaborative sensemaking that emerges in current systems is incidental, not a first-class optimisation target. Richer ecologies offer a complementary lever: multi-agent and open-ended environments show that strategies, tool use, and social conventions emerge when long horizons, other agents, and strategic feedback make them instrumentally valuable baker2020emergent; zhou2024sotopia; zhou2024sotopiapi; park2023generative; qi2024civrealm. To achieve genuine complementarity, we need training ecologies where collaborative friction (disagreement, clarification, and re-framing) can emerge because the environment makes such behaviours rewarding.
Cognitive science shows that humans reason through structured mental models craik1943nature; johnsonlaird1983mental; weick1995sensemaking; klein1998sources, and team effectiveness depends on these models being sufficiently aligned cannonbowers1993shared; mathieu2000influence; mohammed2010teammental. Co-constructing causal structure improves trust and decisions vennix1996group; hovmand2014cbds; constructivist accounts show that learners acquire causal understanding by active exploration, not passive instruction gopnik2004causal; bonawitz2011double. In expert settings there is no single canonical world model available during collaboration, only perspectival models held by particular humans. To be effective, an agent must align with the expert’s causal framing not to blindly validate it, but to obtain a shared reference frame that enables precise error detection and counterfactual critique. This is collaborative causal sensemaking weick1995sensemaking; klein2006dataframe:
Collaborative Causal Sensemaking (CCS): The joint construction, critique, and revision of shared causal and goal models between human and AI, where the agent builds models of how particular experts reason and learns from the outcomes of joint decisions.
Throughout, we illustrate CCS with Lev, an agent assisting physics teacher Dr. Di, whose constructivist philosophy prioritises discovery over drill, and student Ty. When Ty misapplies Newton’s third law, a standard agent prescribes a worksheet; Lev instead surfaces a hypothesis to Dr. Di (“Her diagrams are correct. Could the confusion be linguistic?”), they test it together, and both update their model of Ty’s understanding, which becomes part of a persistent, shared representation that guides future decisions. For Dr. Di, the value is that Lev notices patterns across her students that she would miss, proposes hypotheses grounded in her teaching philosophy, and remembers their joint decisions across weeks, reducing her verification burden while respecting her expertise. A sufficiently prompted LLM can mimic such interactions, but CCS asks: how do we train agents to reliably produce such behaviour, persist the resulting models across sessions, and evaluate whether human–AI alignment actually improves over time?
We propose Collaborative Causal Sensemaking (CCS) as a research agenda for human–AI teams in MAS. Rather than treating collaboration as an interface layer, we argue for training regimes that make collaborative behaviour instrumentally useful: moving from static corpora toward constructivist collaborative playworlds where humans and agents jointly explore, test, and revise explicit causal and goal models to achieve long-horizon objectives baker2020emergent; zhou2024sotopia; park2023generative. In these environments, agents are rewarded not only for task success, but also for sustaining a chain of sensemaking: surfacing hypotheses, generating counterfactual forecasts, and aligning beliefs and priorities over time. CCS spans agent design, reward shaping, and interaction structure to support sensemaking as a capability, grounded in epistemic and teleological alignment clark2025epistemic; li2025hacrl, not just output quality. Our aim is to sharpen what future agents should optimise for in human collaboration.
This framing raises key research questions:
  • Training: What regimes and environment designs elicit collaborative sensemaking rather than polished dialogue?
  • Measurement: How can we formalise and measure alignment (via forecasts, counterfactuals, or causal graphs) without simply rewarding agreement?
  • Transfer: Do CCS behaviours learned in playworlds transfer to deployed decision support (e.g., in shadow-mode deployment and naturalistic logging) so that agents spontaneously initiate sensemaking loops without bespoke prompt engineering?
  • Safety: How can epistemic alignment be operationalised without encouraging agents to manipulate human beliefs?
  • Integration: What bridges are needed between human–AI collaboration research, cognitive science, and large-scale training so that theories of sensemaking shape future MAS pipelines?
Addressing these questions is a precondition for MAS in which agents do not merely answer questions, but think with their human collaborators over time.
2. Agent-Theoretic View of CCS
Before formalising, consider what a CCS agent must track that current agents ignore: the human’s evolving causal beliefs about the domain, their shifting priorities, and the history of where human and agent models have diverged or converged. Standard RL optimises task reward; CCS adds alignment terms that reward shared understanding. We sketch (not prescribe) one way to capture these requirements, using cooperative decision process notation as a scaffold.
From Solitary to Collaborative Decision Processes. We cast expert–assistant interaction as a cooperative, partially observable decision process in the spirit of Dec-POMDPs and cooperative POMDPs oliehoek2016decpomdp; li2025hacrl. At each time $t$, an environment with latent state $s_{t}\in\mathcal{S}$ produces observation $o_{t}\in\mathcal{O}$ (e.g., Ty’s actual understanding of Newton’s third law, observed via quizzes and Dr. Di’s notes) to a human expert $H$ and an assistant $A$. The expert takes actions $a^{H}_{t}\in\mathcal{A}^{H}$ (e.g., Socratic questions, pacing, activity selection), while the assistant takes actions $a^{A}_{t}\in\mathcal{A}^{A}$ (e.g., surfacing diagnostic hypotheses, proposing interventions aligned with Dr. Di’s philosophy). The environment transitions via unknown dynamics $p(s_{t+1}\mid s_{t},a^{H}_{t},a^{A}_{t})$ and yields task rewards $r_{t}$ that both agents ultimately care about.
Crucially, both expert and assistant act through latent world models and goals. We denote by $W^{H}_{t}$ and $W^{A}_{t}$ the internal world models: structured beliefs about task-relevant entities and mechanisms (e.g., Dr. Di’s beliefs about how Ty learns; Lev’s inferences about both). We denote by $G^{H}_{t}$ and $G^{A}_{t}$ their goal structures: representations of what outcomes matter and which objectives should be prioritised (e.g., Dr. Di’s shifting priorities between “build intuition” vs. “cover momentum by Friday”). Both $W_{t}$ and $G_{t}$ evolve as new evidence arrives; they are not fixed exogenous inputs.
In CCS, the relevant system is the team policy $\pi_{I}(a^{H}_{t},a^{A}_{t}\mid\text{history})$ and its joint evolution with $(W^{H}_{t},W^{A}_{t},G^{H}_{t},G^{A}_{t})$. The central question: how to design objectives, data, and architectures that achieve high task return and model convergence.
Alignment Beyond Task Reward: Epistemic and Goal Alignment. We use epistemic alignment to denote alignment in world models and teleological alignment to denote alignment in goals. At a high level, we can think of divergences $d_{W}(W^{A}_{t},W^{H}_{t})$ and $d_{G}(G^{A}_{t},G^{H}_{t})$ that quantify misalignment in causal structure and in objective structure, respectively. If Lev attributes Ty’s error to a diagram gap while Dr. Di suspects a linguistic confusion, $d_{W}>0$; if Lev optimises test scores while Dr. Di prioritises conceptual depth, $d_{G}>0$.
In practice, CCS does not require tracking a full theory-of-mind distribution over an expert’s entire world model or values. A more realistic operating point is local alignment: focusing on the subset of entities, mechanisms, and goals that are currently active in the joint task and aligning those. Factorised or local-graph approximations, where an assistant maintains and revises small, task-specific submodels rather than a monolithic $W^{H}$ and $G^{H}$, offer a plausible route to making CCS-style alignment partially tractable.
In an idealised setting where $W^{H}_{t}$ and $G^{H}_{t}$ were observable, a CCS-style objective might schematically balance task performance with these divergences: $J_{\mathrm{CCS}}\approx\mathbb{E}[\sum_{t}\gamma^{t}r_{t}]-\lambda_{W}\mathbb{E}[d_{W}]-\lambda_{G}\mathbb{E}[d_{G}]$. In practice, the assistant must infer these from actions, language, and co-authored artefacts; $d_{W}$ and $d_{G}$ are instantiated as behavioural proxies over externalised representations (causal sketches, goal descriptions). Moreover, CCS does not demand simple copying: beneficial disagreement requires the assistant to maintain its own hypotheses and surface discrepancies when inferences conflict. In particular, CCS does not advocate aligning to a human’s beliefs regardless of their accuracy: when the assistant’s own model and evidence strongly contradict the expert’s current framing, CCS calls for respectful contestation (surfacing the discrepancy, presenting counter-evidence or alternative causal stories, and supporting the expert in revising their model when appropriate), rather than collapsing to sycophantic agreement.
This sketch connects naturally to existing MAS formalisms. CIRL hadfieldmenell2016cirl treats human–AI interaction as a cooperative game with unknown rewards; CCS extends this to co-evolving world models and goals, not just fixed $\theta$. Active Inference decomposes expected utility into epistemic and pragmatic value friston2024mentalmodels, providing a principled way to trade off information gain about $W$ and $G$ against immediate reward.
The Sensemaking Loop: Discrepancy, Repair, Action. Operationally, CCS manifests as a recurring chain of sensemaking: a loop in which discrepancies between expectations and outcomes trigger collaborative updates to $(W_{t},G_{t})$, followed by revised action. Concretely: (i) Lev notices Ty’s diagrams are correct but verbal explanations wrong (a discrepancy); (ii) Lev and Dr. Di jointly hypothesise a linguistic confusion and test it; (iii) Dr. Di shifts from “re-teach diagrams” to “clarify vocabulary”; (iv) Lev proposes interventions aligned with Dr. Di’s philosophy; she selects and refines weick1995sensemaking; vennix1996group. In human teams, such loops are supported by explicit artefacts (causal maps, after-action reviews, protocols). For CCS in MAS, the research agenda is to design objectives, data, environments, architectures, and interaction policies that make this chain instrumentally valuable for LLM-based agents, so that when deployed, agents default to following such sensemaking loops under naturalistic interaction, rather than requiring bespoke prompting or hand-crafted scripts.
3. Research Agenda for CCS in MAS
Realising CCS in practice requires advances across theory, measurement, data, architectures, and interaction policies. We highlight five intertwined research challenges that map the informal CCS picture into concrete MAS work.
Agenda 1: Formalising Co-Evolving World and Goal Models. Dec-POMDPs, CIRL, and related cooperative frameworks oliehoek2016decpomdp; hadfieldmenell2016cirl; li2025hacrl provide powerful tools for modelling human–AI teams, but they typically assume fixed reward functions, externally specified goals, and do not represent the human’s evolving world model explicitly. CCS instead centres the joint evolution of $(W^{H}_{t},W^{A}_{t},G^{H}_{t},G^{A}_{t})$ as first-class state. We lack MAS formalisms that can represent (i) underdetermined world models that produce identical behaviour on finite data casper2023openproblems, (ii) endogenous goal formation where goals change in response to sensemaking aha2018goalreasoning, and (iii) explicit epistemic and teleological alignment terms as in (2) without collapsing into trivial agreement. Future Directions. Cooperative POMDPs, CIRL, and Active Inference offer ingredients (joint policies, human-aware objectives, epistemic/pragmatic value decompositions friston2024mentalmodels) but none directly represent co-evolving shared world and goal models. Extending these frameworks to include latent $W_{t}$ and $G_{t}$ as state, with dynamics capturing endogenous goal changes (e.g., Dr. Di shifting from “cover momentum” to “repair Newton’s third law” after a surprising quiz), is a key task. Approximations should operate on task-specific abstractions (subgraphs of causal models, goal hierarchy fragments) rather than requiring full theory of mind. Another direction: divergence measures $d_{W}$ and $d_{G}$ compatible with learning, plus regularisers rewarding productive divergence where disagreement triggers tests neither agent would run alone. Formal models of teleological reasoning (inferring latent goals $g_{t}$ that rationalise human actions given $W^{H}_{t}$, as in inverse planning) could be integrated with CCS objectives to ground teleological alignment in observable behaviour.
Agenda 2: Measuring Shared Understanding Without Direct Access. CCS posits that improving epistemic and teleological alignment will reduce verification burden, improve trust calibration, and increase robustness. However, $W^{H}_{t}$ and $G^{H}_{t}$ are latent; we cannot directly compute $d_{W}(W^{A}_{t},W^{H}_{t})$ or $d_{G}(G^{A}_{t},G^{H}_{t})$. Standard metrics for assistants (accuracy, user satisfaction, perplexity) say little about whether human and agent share a compatible causal understanding or goal structure hemmer2024complementarity; clark2025epistemic. An agent may be locally accurate while relying on brittle, spurious patterns; such epistemia (an illusion of knowledge from surface-level associations) is precisely what CCS aims to avoid. Future Directions. A central challenge is to define behavioural and artefact-level proxies for world-model and goal alignment (e.g., do Lev and Dr. Di agree on which students are at risk? how often does Dr. Di override Lev’s suggestions?) and then validate that these proxies are causally linked to collaboration outcomes. When both parties externalise their models as causal graphs, graph-based metrics (e.g., Structural Hamming Distance, graph edit distance) can measure alignment vennix1996group. Counterfactual simulatability tasks test whether human and agent can predict each other’s responses to “what-if” scenarios and future interventions. Team-level evaluation should include verification cost (time and cognitive load spent checking and correcting the assistant), robustness under distribution shift, and complementarity metrics (whether the team outperforms the best individual). Sycophancy stress tests probe whether agents maintain justified beliefs when experts express incorrect opinions—if Dr. Di says “Ty just needs more practice,” does Lev challenge or capitulate? perez-etal-2023-discovering; sharma2024towards Ultimately, we need experimental designs that manipulate alignment (e.g., by perturbing shared models) and test whether this causally affects trust and performance, using proxies informative enough to guide learning yet cheap to elicit.
Agenda 3: Training Ecologies That Reward Sensemaking. Current training corpora consist of static prompt–response pairs, short dialogues, and expert demonstrations ouyang2022instructgpt; zhang2023reasoning. They capture what experts say and do, but not how their $W^{H}_{t}$ and $G^{H}_{t}$ change through discrepancy-driven sensemaking. As a result, agents learn to imitate surface-level behaviour rather than participate in the chain of sensemaking: joint discrepancy detection, causal explanation, goal refinement, and robust action. Future Directions. CCS calls for richer sensemaking trajectories: for Ty, this means (context: repeated errors; anomaly: correct diagrams, wrong explanations; hypotheses: linguistic vs. procedural; disagreement: Lev vs. Dr. Di; repair: vocabulary clarification; goal shift: deprioritise diagrams). Annotation schemes should distinguish epistemic actions (Lev surfacing a hypothesis, Dr. Di probing Ty’s vocabulary) from instrumental actions (executing a chosen plan) lin2025intent. Interactive fine-tuning protocols can log not only corrections but why the expert thinks the assistant erred and how the expert’s own model changed. Naturalistic logging (with governance) can capture genuine goal evolution.
Rather than generic multi-agent simulations, CCS points to constructivist collaborative playworlds engineered as “discrepancy engines”: environments that induce epistemic friction by giving agents partial, biased views of a shared process baker2020emergent; zhou2024sotopia; zhou2024sotopiapi; park2023generative; qi2024civrealm. Such playworlds annotate epistemic moves (noticing mismatches, proposing causal links, renegotiating goals), turning sensemaking trajectories into supervision signals. In such playworlds (e.g., Lev sees quiz logs; Dr. Di sees classroom behaviour; success requires negotiating a shared diagnosis of Ty’s confusion), synthetic experts and assistants can be endowed with different $W$ and $G$ and must align them over time to succeed. Playworlds should be organised into curricula: early levels exercise single-student, single-concept scenarios; later levels involve multi-student patterns, multi-week arcs, and stakeholder conflicts (Dr. Di wants depth; the principal wants test-score gains). Such curricula provide a concrete experimental path: they allow us to study when agents learn clarification, reframing, or goal renegotiation rather than one-shot prediction.
Agenda 4: Architectures for Persistent, Structured Models. LLM-based agents are typically stateless beyond short context windows. They lack persistent, structured world models $W^{A}_{t}$ that can be maintained across tasks, explicit representations of goals $G^{A}_{t}$ that can be revised, and memory systems that record when and why these structures changed. As a result, an agent may learn something important in one interaction and contradict it in the next, or treat transient objectives as if they were stable values. Future Directions. CCS suggests architectural desiderata rather than a single blueprint. Neuro-symbolic causal twins maintain explicit, editable models of the domain that both human and AI can inspect and revise (e.g., a graph where Lev and Dr. Di inspect and edit nodes for Ty, Newton’s laws, and Dr. Di’s goals; when Lev surfaces “Ty: confusion likely linguistic,” Dr. Di can confirm or reject), serving as shared artefacts for sensemaking vennix1996group; grieves2017digitaltwin. In such architectures, LLMs serve as flexible “epistemic encoders” that translate language and observations into edits on an explicit causal and goal model, while a lightweight reasoner checks consistency, supports counterfactual prediction, and records provenance.
Episodic sensemaking memory stores triplets such as (Ty answered diagrams correctly; verbal explanations wrong; Dr. Di deprioritised diagram remediation), enabling Lev to learn: “correct-procedure-wrong-explanation triggers vocabulary investigation for Dr. Di.” Teleological representations such as reward machines icarte2018rewardmachines can encode the logical structure of goals; joint inference over these machines and causal graphs can link epistemic updates (editing $W$) to teleological updates (editing $G$). A lightweight theory-of-mind module maintains hypotheses about $W^{H}_{t}$ and $G^{H}_{t}$; Lev models Dr. Di’s preference for discovery, framing recommendations accordingly. In deployment, whether CCS behaviours actually manifest will depend on how these structures are threaded through context windows: episodic memories must survive across interactions and be retrieved at the right time to automatically shape future recommendations.
Agenda 5: When to Disagree, When to Defer. Even with appropriate objectives, data, and architectures, we lack principled policies for when CCS agents should agree, challenge, ask clarifying questions, or slow interaction for epistemic repair Bansal2021UpdatesHarm; alonbarkat2023trust; li2025hacrl. Current assistants are optimised for low-friction helpfulness: they answer quickly, avoid conflict, and rarely question the user’s framing. Effective collaborators must sometimes do the opposite: pause, surface uncertainty, or propose goal revisions. At the same time, CCS introduces new risks: agents that infer and update goals endogenously may develop goal structures that drift away from human intent; agents trained to avoid sycophancy may become overconfident or manipulative. Future Directions. Beyond what to say, CCS raises questions about when an agent should surface discrepancies and slow interaction for epistemic repair instead of answering fluently and moving on. Value-of-Information criteria lu2024activereThinking can estimate an expected benefit of repair, trading off uncertainty reduction, outcome criticality, and friction cost (e.g., should Lev interrupt to flag Ty’s risk, or trust Dr. Di’s long-game pedagogy?). Mixed-initiative protocols can formalise turn-taking and control: when the assistant is allowed to override, when it must defer, and when it suggests after-action reviews. Training for “intelligent disobedience” can teach agents to contest risky decisions. If Dr. Di says “just give Ty the worksheet,” Lev might respond: “The worksheet raises scores, but the confusion may resurface in momentum. Flag for review?”
CCS systems will need teleological constraints: constitutional principles that bound goal formation and prevent agents from extrapolating goals in undesirable ways. Avoiding both sycophancy and “sycophancy inversion” requires adaptive personalisation based on expertise, context, and stakes. High-stakes sensemaking should be auditable via epistemic provenance trails (e.g., “Dr. Di rejected three drill recommendations; inferred: prefers discovery”; “Dr. Di said use swimming for Ty; inferred: contextualise to interests”) hao2025beyondhitl. These concerns connect CCS to broader debates on accountability and human-in-the-loop oversight in MAS.
4. Conclusion
We have argued that making LLM-based agents into genuine teammates in MAS for decision support requires shifting from behavioural alignment to collaborative causal sensemaking: the joint construction, critique, and revision of shared world and goal models that underpin decisions. Rather than treating collaboration as an interface layer, CCS treats the human’s evolving mental models and objectives as part of the decision state that agents must track, stress-test, and help refine. We sketched an agent-theoretic view in which epistemic and teleological alignment appear alongside task reward, and outlined research challenges in formalisation, measurement, playworld design, architectures, and interaction policies. The central hypothesis is that such alignment can reduce verification burden while enabling calibrated reliance and productive disagreement, with near-term footholds in CCS playworlds, causal-twin prototypes, and shadow-mode deployment where agents must demonstrate these behaviours under naturalistic conditions. Where instruction tuning builds tools that obey, CCS aims to build teammates that participate in the reasoning behind choices and think with their human partners.